X + 5 + X + X - 45 = 392
I come up with X = 144?

Found 2 solutions by Fombitz, Nieleya:
Answer by Fombitz(32388)

Yes, you're correct!

Answer by Nieleya(22)

X + 5 + X + X - 45 = 392
I come up with X = 144?
3x + 5 - 45 = 392
3x - 40 = 392
3x = 392 + 40
3x = 432
x = 432/3
x = 144